OpenStudy (calculusxy):

Find the sum of the arithmetic series. 1 + 4 + 7 + 10 + ... + 43

OpenStudy (calculusxy):

@ganeshie8

OpenStudy (samsungfanboy):

how??? ;-;

rebeccaxhawaii (rebeccaxhawaii):

okay well a_1=1

OpenStudy (calculusxy):

yes

rebeccaxhawaii (rebeccaxhawaii):

how much is it increasing each time?

OpenStudy (calculusxy):

3

rebeccaxhawaii (rebeccaxhawaii):

correct so d is 3

OpenStudy (calculusxy):

I am not sure, but does this work to present this situation? \[\sum_{n=1}^{43}n + 3\]

rebeccaxhawaii (rebeccaxhawaii):

to get the sum this is your formula \[s_n=\frac{ n }{ 2 }[a_1+a_n]\]

OpenStudy (calculusxy):

yes

rebeccaxhawaii (rebeccaxhawaii):

plug in a_1=1 and a_n=43 and n=?

rebeccaxhawaii (rebeccaxhawaii):

but first we need to find n

rebeccaxhawaii (rebeccaxhawaii):

which n=14

OpenStudy (calculusxy):

how did you get 14?

rebeccaxhawaii (rebeccaxhawaii):

kept adding 3 in my head till i got 43

OpenStudy (calculusxy):

ok. is there any algebraic way to find n?

OpenStudy (calculusxy):

i mean i know that if we know what \(S_n\) then it can be possible, but in this case would it work?

rebeccaxhawaii (rebeccaxhawaii):

yes lemme get it for you

OpenStudy (calculusxy):

ok

rebeccaxhawaii (rebeccaxhawaii):

1+(n-1)*3=43 n=15 not 14 sorry

OpenStudy (calculusxy):

okay so you use the explicit formula to find this right?

rebeccaxhawaii (rebeccaxhawaii):

a_1+(n-1)*d=a_n yes this is explicit

OpenStudy (calculusxy):

thanks! can you help me with one more problem?

rebeccaxhawaii (rebeccaxhawaii):

I can sure try to! Tag me in your new thread.